OpenStudy (anonymous):

how do you factor g^2-5g-14 ?

OpenStudy (anonymous):

do you know 14 is multiple by what number?

OpenStudy (anonymous):

(x+a)(x+b)=x^2+(a+b)x+ab you must find a and b that ab=-14 (a+b)=-5